Nuclear power stations in Canada use a type of reactor inventea in Can a Canada Deuterium Uranium (CANDU) reactor. A typical CANDU fuel bu mass of 23 kg. A A mada called bundle has a (a) Calculate the amount of electrical energy produced by a single CANDU fuel bundh assuming that the entire bundle is converted to electrical energy. [ans:2.1x 10 undle (b) An average Canadian home uses 3.6 x 1010 J of energy per day. Determine the number of days that a single CANDU fuel bundle could theoretically provi energy to a home. [ans: 5.8 x 10 days]
Answer #1

Q1. (a)

Ans- Amount of electrical energy produced by a single CANDU fuel bundle is,

E=mc^2 =23 * (3*10^8)^2 = 2.1 * 10^18 J

So,  Amount of electrical energy produced by a single CANDU fuel bundle is  2.1 * 10^18 J.

(b)

Ans- enegy used by a Canadian home per day is , E1= 3.6 * 10^10 J

So, Number of days a CANDU fuel bundle can provide energy to a home is

= E/E1 =  (2.1 * 10^18 J )/(3.6 * 10^10J) =5.8 * 10^7 days

So, Number of days a CANDU fuel bundle can provide energy to a home is 5.8 * 10^7 days